### v2

# Improving Model’s Interpretability and Reliability using Biomarkers

Abstract
Accurate and interpretable diagnostic models are crucial in the safety-critical field of medicine. We investigate the interpretability of our proposed biomarker-based lung ultrasound diagnostic pipeline to enhance clinicians’ diagnostic capabilities. The objective of this study is to assess whether explanations from a decision tree classifier, utilizing biomarkers, can improve users’ ability to identify inaccurate model predictions compared to conventional saliency maps. Our findings demonstrate that decision tree explanations, based on clinically established biomarkers, can assist clinicians in detecting false positives, thus improving the reliability of diagnostic models in medicine.
1 Introduction
The development of accurate and dependable diagnostic models holds significant importance in the safety-critical field of medicine. Our research, described in [4], proposes a learning mechanism where the neural network model is enforced to go through an interpretable feature bottleneck of clinically established lung ultrasound biomarkers. As the feature encoder extracts biomarkers well-known to ultrasound radiologists, such expert users can now easily verify the black-box feature extractor output. The downstream classifiers (such as decision trees) that operate on these biomarkers are inherently interpretable as they operate on known features. Consequently, we have achieved a fully interpretable diagnostic model that adheres to the interpretability principles outlined in [7]. In this work, we conduct a user study to formally assess the interpretability of our proposed biomarker model.
Figure 1: In contrast to conventional approach that learn task-specific feature, we proposed in [4] the decoupling of feature extraction from end-tasks by enforcing models to go through an interpretable feature bottleneck of clinically established biomarkers. With help of clinicians we defined 38 lung ultrasound biomarkers that match/exceed DNN performance especially at low data regime. As the feature encoder extracts biomarkers well-known to ultrasound radiologists, clinician can now easily verify the black-box feature extractor output. Also, the downstream classifiers (such as decision trees) that operate on these biomarkers are diagnosable as they operate on known features. Thus giving rise to a fully interpretable diagnostic model.
Lung ultrasound has emerged as a dependable diagnostic tool for clinicians, particularly in the context of the COVID-19 pandemic [2]. This has spurred a rapid development of AI-based approaches for interpreting lung ultrasound images to aid in disease diagnosis [8, 1, 3, 10, 5, 6]. As the use of AI-based diagnostic models becomes more widespread, there is a pressing need for reliable and interpretable techniques. These techniques are essential to ensure that clinicians can gain a clear understanding of the workings of these ”black-box” models. This understanding is crucial for accurate and safe patient care, particularly in the safety critical field of medicine.
The objective of this study is to evaluate whether explanations obtained from a decision tree classifier (built on our biomarkers [4]) improves the ability of users to identify inaccurate model predictions. Decision explanations depict the decision logic that the model used to derive its predictions. Saliency maps are a widely used post-hoc analysis tool that provides valuable information by highlighting the input image regions that influence the classifier’s output [9]. However, saliency maps cannot quantify the effect magnitude nor effect direction of the observed area of focus, limiting its utility in providing insight into the model’s decision-making. The use of both saliency maps and decision trees may provide additional insights into the decision process and the evidence utilized by the model to make its predictions. Using this added insight, users may be able to more accurately determine when a model’s output should be accepted and when it should be rejected.
Figure 2: Diagnoising the model’s mis-prediction of lung-severity score as 0 instead of 1 for the above ultrasound clip, by examining the decision-tree explanations, we can easily discern that the model overlooked the presence of B-lines, leading to the misprediction. In contrast, relying solely on the saliency map makes it challenging to identify the cause of the misprediction, as the model’s attention is distributed across various regions in the frames, including the B-line. Consequently, making it difficult to quantify the contribution of the B-line to the final prediction.
2 Methods
We conduct the study on 30 lung ultrasound (LUS) patient videos from the test dataset of [4]. Of these videos, the model accurately predicted the lung-severity score (0, 1, 2, or 3) in 18 cases (60% accuracy). Using the MLP classifier, we generated Grad-CAM [9] based saliency maps, and we extracted the decision path of the videos from the decision tree classifier. Both the MLP and the decision tree classifiers produced identical predictions for all videos.
For each video clip, the clinicians were presented with the model’s predicted lung severity score and were asked to evaluate its correctness with the assistance of either: saliency map, decision tree explanation, both, or none. The presentation of the videos and their variants were randomized to ensure that each evaluation was independent. The study was conducted with three LUS clinicians.
Figure 3: Box plot showing the clinician’s detection rates of the correctness of AI’s output. We see that decision trees are effective at helping users detect false positive predictions. Overall we see that it’s beneficial to have a model analysis tool to improve over baseline accuracy of 60%, with saliency maps being the most beneficial.
3 Results
Figure 3 shows that saliency maps are good at detecting true positives, whereas the decision tree is good at detecting false positives. Presenting both results in having the lowest standard deviation, indicating greater consensus among the clinicians. Finally, after going through all 30 clips, they rated the usefulness of the decision trees as 0.51 and heat-maps as 0.02 (average score on a scale of -1 to 1) and two clinicians favored having both while the third preferred decision trees over saliency maps.
Figure 4: With our decision-tree explanations we can easily interpret the video. Which would be helpful for novice clinicians, who in this case could mistake the vertically stacked A-line bands for B-lines, whereas our decision tree explanation clearly shows no B-lines are present.
Discussion: We find that interpreting ultrasound videos is more straightforward using our decision-tree explanations. This feature is particularly beneficial for novice clinicians. For example, in the ultrasound clip shown in Figure 4, a novice clinician might mistakenly identify the vertically stacked A-line bands as B-lines. However, our decision tree explanation clearly indicates the absence of B-lines.
In contrast, using saliency maps for interpreting ultrasound videos could be misleading. For example, in the ultrasound clip depicted in Figure 2, the model mispredicted the lung severity score as 0 instead of 1. Relying solely on the saliency map to diagnose this misprediction makes it challenging to identify the cause, as the model’s attention is distributed across various regions in the frames, including the B-line. Consequently, quantifying the contribution of the B-line to the final prediction becomes difficult. However, with the decision-tree explanations, we can easily determine that the model overlooked the presence of B-lines, which led to the misprediction.
4 Conclusion
We conducted a user study comparing the interpretability of our proposed biomarker based lung ultrasound diagnostic pipeline. We compared decision tree explanations built on top of biomarkers with conventional saliency maps to help clinicians detect model mis-predictions. We see that access to analysis tools helps improve the reliability of models, where decision tree classifiers built on top of our biomarkers can help detect false positives, vital for diagnostic medicine.
Acknowledgments: This research is supported in part by the Center for Machine Learning and Health (CMLH) Translational Fellowship, Carnegie Mellon University (CMU). The authors thank the clinicians at Louisiana State University (LSU) for their assistance with data collection and our CMU collaborators for their valuable insights. All data used in this study were collected under IRB protocol number 1509, titled Artificial Intelligence Interpretation of Lung Ultrasound Images, and were de-identified prior to transfer to CMU.